OpenStudy (anonymous):

Find the center, vertices, and foci of the ellipse with equation 2x2 + 7y2 = 14.

OpenStudy (anonymous):

divide by 14 \[\frac{ \left( x-h \right)^{2} }{ a ^{2} }+\frac{ \left( y-k \right)^{2} }{ b ^{2} }=1\] centre is (h,k) \[b ^{2}=a ^{2}\left( 1-e ^{2} \right)\] foci are (-ae,k) and (ae,k)

OpenStudy (anonymous):

um to be honest I am not even really sure how to plug my equation into that @surjithayer

OpenStudy (anonymous):

this is really confusing...

OpenStudy (anonymous):

actually foci are (h-ae,k) and (h+ae,k) vertices are (h-a,k) and (h+a,k)

OpenStudy (anonymous):

let us start divide by14

OpenStudy (anonymous):

\[\frac{ x ^{2} }{7 }+\frac{ y ^{2} }{2 }=1\]

OpenStudy (anonymous):

here centre is (0,0)

OpenStudy (anonymous):

\[a=\sqrt{7},b=\sqrt{2}\]

OpenStudy (anonymous):

now you calculate e

OpenStudy (anonymous):

\[2=7\left( 1-e ^{2} \right),\frac{ 2 }{7 }=1-e ^{2},e ^{2}=1-\frac{ 2 }{ 7 },e=\sqrt{\frac{ 5 }{ 7 }}\]

OpenStudy (anonymous):

foci are (-ae,0) and (ae,0) vertices are (-a,0) and (a,0)